OpenStudy (anonymous):

Solve 5(4x-7) + 12 = 17

OpenStudy (anonymous):

x = 2

OpenStudy (anonymous):

20x-35+12=17 20x-23 = 17 20x= 17+23 x = 40/20 x = 2